OpenStudy (anonymous):

I need help... In school we just learned how to add fractions together with unlike denominators... Dont really remember could someone help please

OpenStudy (amistre64):

I would simply equate them to some unknown number: 1/2+1/4.= N ; and let the math solve itself 1(2)/2+1(2)/4.= (2)N 1+1/2.= 2N 1(2)+1(2)/2.= 2(2)N 2+1= 4N 3= 4N 3/4= 4N/4 3/4= N

OpenStudy (anonymous):

1/4 + 2/3?

OpenStudy (amistre64):

same steps; the process doesnt change; only the numbers that are used

OpenStudy (anonymous):

(1/4)*3 + (2/3)*4 = 3/12+8/12 = 11/12

OpenStudy (amistre64):

1/4 + 2/3 = N 1(3)/4 + 2(3)/3 = (3)N 3/4 + 2 = 3N 3(4)/4 + 2(4) = 3(4)N 3 + 8 = 12N 11 = 12N 11/12 = 12N/12 11/12 = N

OpenStudy (anonymous):

Thankz ppl!

OpenStudy (amistre64):

yep